An air bubble has a volume of 2.5 cm3 when it is released at a depth of 40 m of water. What will be its volume when it reaches the surface of the water? Assume that its temperature is constant and that atmospheric pressure = 10 m of water.
1
Expert's answer
2019-03-29T11:01:46-0400

We assume that its temperature is constant. Applying Boyle's law:


"PV = constant"

"(P + H\u03c1g)V = P V'"

"V'=V(1+\frac{Hρg}{P})"

We have:


"P=h\u03c1g"

Thus,

 

"V'=V(1+\frac{Hρg}{hρg})=V(1+\frac{H}{h})"

"V'=2.5(1+\frac{40}{10})=12.5 cm^3"